# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v2

@@ 6 Conclusion and Discussion
Our analysis of LLMs through the lens of latent variable models and identifiability provides valuable insights into the underlying mechanisms that contribute to their success [^1] We strongly encourage the reader to refer to the additional discussions on latent variable models in Appendix G.. We have shown that that the representations learned by LLMs can be approximated as a linear transformation of the posterior distribution of latent variables. This not only reinforces the linear representation hypothesis but also highlights that concepts are encoded in the linear transformation matrix.
The fact that the representation learned by LLMs approximates a linear transformation of the underlying causal model of the observed data shows that next token prediction is enough to support learning a deep understanding of the data. Based on this finding, we suggest the following directions:
Embedding Causal Reasoning in LLMs Through Linear Unmixing
Our linear identifiability result lays a foundation for uncovering latent causal relationships among variables, especially when these variables exhibit causal dependencies within the proposed latent variable model. By showing that the representations in LLMs are linear mixtures of latent causal variables, our analysis shows that linear unmixing of these representations may allow for the identification of underlying latent causal structures. This approach not only opens up the possibility of understanding causal dynamics within LLMs but also suggests that causal reasoning, particularly in latent spaces, could be achievable through the exploration of these linear unmixing techniques. We believe this work marks a pivotal step toward embedding robust causal reasoning capabilities into LLMs.
@@ Appendix A Related Work
Recent studies have established the empirical finding that concepts in LLMs are often linearly encoded, a phenomenon known as the linear representation hypothesis (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Park et al., 2023; Li et al., 2024; Gurnee et al., 2023). Building on this observation, recent works (Park et al., 2023; Marconato et al., 2024) attempt to unify these findings into a cohesive framework, aiming to deepen our understanding and potentially inspire new insights. However, these works do not address why and how such linear properties emerge. Some previous works (Marconato et al., 2024; Roeder et al., 2021) have demonstrated identifiability results within the inference space, such as establishing connections between features derived from different inference models. However, these results are confined to the inference space and do not connect to the true latent variables in latent variable models. A recent work (Jiang et al., 2024) seeks to explain the origins of these linear properties but employs a latent variable model that differs from ours. From a technical perspective, their explanation is rooted in the implicit bias of gradient descent. In contrast, our work provides an explanation grounded in identifiability theory. This distinction highlights our focus on connecting the observed linear properties directly to the identifiability of true latent variables in latent variable models, offering a more comprehensive and theoretically robust understanding of these phenomena.
@@
This work is closely related to causal representation learning (Schölkopf et al., 2021), which aims to identify high-level latent causal variables from low-level observational data. Many prior studies (Brehmer et al., 2022; Von Kügelgen et al., 2021; Massidda et al., 2023; von Kügelgen et al., 2023; Ahuja et al., 2023; Seigal et al., 2022; Shen et al., 2022; Liu et al., 2022; Buchholz et al., 2023; Varici et al., 2023; Liu et al., 2024b, 2025, a, c; Hyvarinen & Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020) have developed theoretical frameworks supporting the recovery of true latent variables up to simple transformations. However, these works primarily focus on continuous spaces and do not address the next-token prediction framework employed by LLMs. A subset of works (Gu & Dunson, 2023; Kong et al., 2024; Kivva et al., 2021) has explored causal representation learning in discrete spaces for both latent and observed variables, often imposing specific graph structures and assuming invertible mappings from latent to observed spaces. Despite their focus on discrete settings, none of these studies consider the next-token prediction framework in LLMs. In contrast, our work overcomes these limitations. We analyze approximate identifiability without relying on strict invertibility assumptions, which better aligns with the complex and often non-invertible relationships observed in real-world data. While a recent study has examined non-invertible mappings from latent to observed spaces, they rely on additional historical information to effectively restore invertibility. In our approach, we focus on achieving approximate identifiability without requiring such additional information. A very recent work (Rajendran et al., 2024) explores identifiability analysis for LLMs; however, they model both observed text data and latent variables as continuous and still assume invertibility in their analysis. In contrast, our framework explicitly considers discrete latent and observed variables, relaxed invertibility assumptions, and directly aligns with the next-token prediction paradigm, offering a more realistic and generalizable approach to causal representation learning in LLMs.
@@ Appendix B Limitations
One of the main contributions of this work is establishing an identifiability result for the next-token prediction framework, a widely used approach for training LLMs. This result hinges on a key assumption: the diversity condition, which requires the data distribution to exhibit sufficiently strong and diverse variations. The diversity condition was originally introduced in the context of nonlinear ICA (Hyvarinen & Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020) and has since been adopted in the identifiability analysis within the causal representation learning community. In essence, it has emerged as a fundamental requirement for identifiability. This condition is likely satisfied given that the training data for LLMs is primarily sourced from the Internet, encompassing a broad and diverse range of content. Recent empirical analyses of LLMs further support the plausibility of this assumption (Roeder et al., 2021; Marconato et al., 2024), reinforcing its relevance in the context of identifiability.
Similar to existing works (Park et al., 2023; Jiang et al., 2024; Marconato et al., 2024; Rajendran et al., 2024), our analysis is limited to a specific layer in LLMs and does not provide a justification for intermediate layers. Formalizing the relationship between features in intermediate layers and identifiability analysis presents additional challenges. This complexity arises from the intricate nature of intermediate layer representations. Therefore, extending identifiability results to intermediate layers remains an open and more challenging problem.
@@
Next-token prediction mirrors multinomial logistic regression, where the model approximates the true posterior distribution, i.e., $p(y|\mathbf{x})$, by minimizing the cross-entropy loss. In this framework, given sufficient training data, a sufficiently expressive architecture, and effective optimization, the model’s predictions are expected to converge to the true posterior, as follows:
$p(y|\mathbf{x})=\frac{\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y))}{\sum_{y}\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y))}.$ (6)
On the other hand, for the proposed latent variable model (illustrated on the left of Figure 1 or described in Eq. 1), the posterior distribution $p({y}|\mathbf{x})$ can be derived using Bayes’ rule:
$p({y}|\mathbf{x})=\sum_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}).$ (7)
@@
$\frac{\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y))}{\sum_{y}\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y))}=\sum_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}).$ (8)
Taking the natural logarithm on both sides of Eq. (8), we obtain:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log{Z(\mathbf{x})}=\log\big{(}\sum_{c}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})\big{)}$ (9)
where ${Z(\mathbf{x})}=\sum_{y}\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y))$ represents the normalization constant. This transformation provides a direct link between the posterior approximation derived from next-token prediction (Eq. (6)) and the Bayesian formulation of the latent variable model (Eq. (7)). It enables a detailed comparison of how the latent variable decomposition $p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})$ interacts with the parameterized representation $\mathbf{f}_{x},\mathbf{f}_{y}$ in a log-linear form.
Now, let us focus on the right of Eq. (8). Introducing $p(\mathbf{c}|y)$ and $q(\mathbf{c})$, where $q(\mathbf{c})$ is a valid probability distribution such that $\sum_{c}q(\mathbf{c})=1$ , we can obtain that:
$\log p(y|\mathbf{x})$ $=\mathbb{E}_{p(\mathbf{c}|y)}\log p(y|\mathbf{x})$ (10)
$=\mathbb{E}_{p(\mathbf{c}|y)}\log\big{(}p(\mathbf{c}|y,\mathbf{x})\frac{p(y|\mathbf{x})}{p(\mathbf{c}|y,\mathbf{x})}\big{)}$ (11)
$=\mathbb{E}_{p(\mathbf{c}|y)}\log{p(y,\mathbf{c}|\mathbf{x})}-\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|y,\mathbf{x})}$ (12)
$=\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}+\mathbb{E}_{p(\mathbf{c}|y)}\log{\underbrace{p(y|\mathbf{c},\mathbf{x})}_{p(y|\mathbf{c})}}-\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|y,\mathbf{x})}$ (13)
@@
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log{Z(\mathbf{x})}=\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}+\underbrace{\mathbb{E}_{p(\mathbf{c}|y)}\log{{p(y|\mathbf{c})}}}_{b_{y}}-\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|y,\mathbf{x})}.$ (14)
Define $\mathbb{E}_{p(\mathbf{c}|y)}\log{{p(y|\mathbf{c})}}=b_{y}$ Define a vector $[p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}$ as the vector constructed from the probabilities of all possible values of $\mathbf{c}$, conditional on $\mathbf{x}$, where ${\mathbf{c}_{i}\in\mathcal{C}}$ and $\mathcal{C}$ is the set of all possible values that $\mathbf{c}$ can take. As a result, $\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}=\sum_{\mathbf{c}}{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}=[{p}(\mathbf{c}=\mathbf{c}_{i}|y)]_{\mathbf{c}_{i}}{[\log p(\mathbf{c}|\mathbf{x})]_{\mathbf{c}_{i}}}$. Then we can re-write Eq. (14) as follows:
@@
Now, we apply the diversity condition outlined in Section 3, which asserts that there exist $\ell+1$ values of $y$, where $\ell$ is the number of all possible values of $\mathbf{c}$, i.e., $y_{0}$, $y_{1}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})]^{T}_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]^{T}_{\mathbf{c}_{i}},...,[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})]^{T}_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]^{T}_{\mathbf{c}_{i}}\big{)}$ of size $\ell\times\ell$ is invertible. In this context, for $y=0$, we have:
@@
$\big{(}\mathbf{f}_{y}(y=y_{1})^{T}-\mathbf{f}_{y}(y=y_{0})^{T}\big{)}\mathbf{f}_{x}(\mathbf{x})=$ $\big{(}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})}]^{T}_{\mathbf{c}_{i}}-[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]^{T}_{\mathbf{c}_{i}}\big{)}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$
@@
According to the diversity condition, where $y$ can take $\ell+1$ values, we obtain a total of $\ell$ equations similar to Eq. (19). Collecting all of these equations, we have:
@@
$=\underbrace{\big{(}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})}]_{\mathbf{c}_{i}}-[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]_{\mathbf{c}_{i}},...,[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})}]_{\mathbf{c}_{i}}-[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]_{\mathbf{c}_{i}}\big{)}^{T}}_{\mathbf{L}}\times$
@@
Now, we focus on the two terms in the right of Eq. (22), i.e., $\mathbf{{h}}_{y}$ and $\mathbf{{b}}_{y}$, from the viewpoints of when the mapping from latent space, e.g., $\mathbf{c}$, to observed space, e.g., $\mathbf{x}$ and $y$ is invertible and approximately invertible.
When the mapping $\mathbf{g}$ from latent space to observed space is invertible.
We assume that the mapping from the latent space $\mathbf{c}$ to the observed space $(\mathbf{x},y)$ is approximately invertible, meaning that
@@
where $10\leq\epsilon<$ 1, and $\mathbf{c}^{*}=\arg\max_{\mathbf{c}}p(\mathbf{c}\mid\mathbf{x},y)$ denotes the dominant latent configuration. When the mapping $\mathbf{g}$ is exactly invertible, we have $\epsilon=0$, implying perfect identifiability of the latent variable from the observed data.
  • For the quantity $\mathbf{h}_{y}$, we analyze each component, i.e., $h_{y_{i}}-h_{y_{0}}$, where
@@
    which implies that
    $\log p(\mathbf{c}\mid\mathbf{x},y)=\log\delta(\mathbf{c}-\mathbf{c}^{*}).$ (26)
    In this case, since the posterior concentrates on a single value, its expectation under any distribution $p(\mathbf{c}\mid y)$ becomes constant, and thus
@@
  • Similarly, Recall the definition of the quantity $\mathbf{b}_{y}$. For any pair of labels $y_{i}$ and $y_{0}$, the difference between components is given by:
    $b_{y_{i}}-b_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(y=y_{i}\mid\mathbf{c})\right]-\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(y=y_{0}\mid\mathbf{c})\right].$ (28)
    Now since the mapping from latent variable $\mathbf{c}$ to observed pair $(\mathbf{x},y)$ is invertible, whch implies that for each $y=y_{k}$, the conditional distribution $p(\mathbf{c}\mid y=y_{k})$ is supported only on values of $\mathbf{c}$ such that $p(y=y_{k}\mid\mathbf{c})=1$. That is:
    $p(y=y_{k}\mid\mathbf{c})=1\quad\text{for }\mathbf{c}\sim p(\mathbf{c}\mid y=y_{k}).$ (29)
    Therefore, for any such $\mathbf{c}$, we have:
    $\log p(y=y_{k}\mid\mathbf{c})=\log 1=0.$ (30)
    Hence, each term in the difference becomes:
    $\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\log p(y=y_{i}\mid\mathbf{c})=0,$ (31)
    $\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\log p(y=y_{0}\mid\mathbf{c})=0.$ (32)
    Thus, the difference is:
    $b_{y_{i}}-b_{y_{0}}=0-0=0.$ (33)
When the mapping $\mathbf{g}$ is approximately invertible, i.e., $\epsilon\rightarrow 0$,
  • For the quantity $\mathbf{h}_{y}$, We analyze the component difference:
    $h_{y_{i}}-h_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(\mathbf{c}\mid\mathbf{x},y_{i})\right]-\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(\mathbf{c}\mid\mathbf{x},y_{0})\right].$
@@
    $p(\mathbf{c}\mid\mathbf{x},y)=(1-\epsilon)\cdot\delta_{\mathbf{c}^{*}}+\epsilon\cdot q(\mathbf{c}),$
@@
    $\mathbb{E}_{p(\mathbf{c}\mid y)}\left[\log p(\mathbf{c}\mid\mathbf{x},y)\right]=p(\mathbf{c}^{*}\mid y)\log(1-\epsilon)+\sum_{\mathbf{c}\neq\mathbf{c}^{*}}p(\mathbf{c}\mid y)\log p(\mathbf{c}\mid\mathbf{x},y).$
@@
    $\mathbb{E}_{p(\mathbf{c}\mid y)}\left[\log p(\mathbf{c}\mid\mathbf{x},y)\right]\approx\log\epsilon,$

### v3

@@ 7 Conclusion
Our analysis of large language models (LLMs) through the lens of latent variable models and identifiability offers key insights into the mechanisms driving their success [^2] We encourage the reader to refer to the additional discussions on latent variable models in Appendix I.. We have demonstrated that the representations learned by LLMs can be effectively approximated as a linear transformation of the posterior distribution of latent variables. This not only provides an initial framework for understanding next-token prediction, but also highlights the underlying linear properties of LLMs, including support for the linear representation hypothesis and sparse autoencoders. These findings pave the way for further exploration of how LLMs learn and represent complex patterns in data. Based on this finding, we suggest the following directions: 1) Rethinking Invertibility Assumptions in Causal Representation Learning, 2) Embedding Causal Reasoning in LLMs Through Linear Unmixing. Refer to Appendix K for more details.
@@ Appendix A Related Work
Recent studies have established the empirical finding that concepts in LLMs are often linearly encoded, a phenomenon known as the linear representation hypothesis [54, 60, 5, 23, 16, 69, 57, 56, 59, 43, 31, 63]. Building on this observation, recent works [59, 51] attempt to unify these findings into a cohesive framework, aiming to deepen our understanding and potentially inspire new insights. However, these works do not address why and how such linear properties emerge. Some previous works [51, 64] have demonstrated identifiability results within the inference space, such as establishing connections between features derived from different inference models. However, these results are confined to the inference space and do not connect to the true latent variables in latent variable models. A recent work [35] seeks to explain the origins of these linear properties but employs a latent variable model that differs from ours. From a technical perspective, their explanation is rooted in the implicit bias of gradient descent. In contrast, our work provides an explanation grounded in identifiability theory. This distinction highlights our focus on connecting the observed linear properties directly to the identifiability of true latent variables in latent variable models, offering a more comprehensive and theoretically robust understanding of these phenomena. In addition, recent work [63] assumes continuous latent and observed variables, while we assume both latent variables and observed variables to be discrete, aligning more closely with modeling natural language.
@@
This work is closely related to causal representation learning [66], which aims to identify high-level latent causal variables from low-level observational data. Many prior studies [12, 74, 53, 75, 3, 67, 68, 44, 15, 73, 46, 48, 45, 47, 33, 34, 39] have developed theoretical frameworks supporting the recovery of true latent variables up to simple transformations. However, these works primarily focus on continuous spaces and do not address the next-token prediction framework employed by LLMs. A subset of works [29, 41, 40] has explored causal representation learning in discrete spaces for both latent and observed variables, often imposing specific graph structures and assuming invertible mappings from latent to observed spaces. Despite their focus on discrete settings, none of these studies consider the next-token prediction framework in LLMs. In contrast, our work overcomes these limitations. We analyze approximate identifiability without relying on strict invertibility assumptions, which better aligns with the complex and often non-invertible relationships observed in real-world data. While a recent study has examined non-invertible mappings from latent to observed spaces, they rely on additional historical information to effectively restore invertibility. In our approach, we focus on achieving approximate identifiability without requiring such additional information. A very recent work [63] explores identifiability analysis for LLMs; however, they model both observed text data and latent variables as continuous and still assume invertibility in their analysis. In contrast, our framework explicitly considers discrete latent and observed variables, relaxed invertibility assumptions, and directly aligns with the next-token prediction paradigm, offering a more realistic and generalizable approach to causal representation learning in LLMs.
Sparse Autoencoders
Polysemanticity, a phenomenon observed in recent studies, roughly speaking, refers to cases where a single representation encodes multiple distinct, human-interpretable concepts [6]. Early investigations suggested that neural networks represent features by linear superposition and motivated efforts to disentangle human-interpretable concepts from such linear mixing [23]. This can be achieved by using sparse autoencoders [32, 61, 62, 27, 11, 13], a technique closely related to the well-known framework of dictionary learning [19, 20, 21, 22, 2, 4]. In contrast to these works, we propose structured SAEs to model the dependencies among latent concepts. Furthermore, motivated by our theoretical findings, we introduce a new evaluation method for SAEs, grounded in our justified theoretical results.
@@ Appendix B Limitations
One of the main contributions of this work is establishing an identifiability result for the next-token prediction framework, a widely used approach for training LLMs. This result hinges on a key assumption: the diversity condition, which requires the data distribution to exhibit sufficiently strong and diverse variations. The diversity condition was originally introduced in the context of nonlinear ICA [33, 34, 39] and has since been adopted in the identifiability analysis within the causal representation learning community. In essence, it has emerged as a fundamental requirement for identifiability. This condition is likely satisfied given that the training data for LLMs is primarily sourced from the Internet, encompassing a broad and diverse range of content. Recent empirical analyses of LLMs further support the plausibility of this assumption [64, 51], reinforcing its relevance in the context of identifiability.
Similar to existing works [59, 35, 51, 63], our analysis is limited to a specific layer in LLMs and does not provide a justification for intermediate layers. Formalizing the relationship between features in intermediate layers and identifiability analysis presents additional challenges. This complexity arises from the intricate nature of intermediate layer representations. Therefore, extending identifiability results to intermediate layers remains an open and more challenging problem.
@@
Proof.
Next-token prediction mirrors multinomial logistic regression, where the model approximates the true conditional distribution, i.e., $p(y|\mathbf{x})$, by minimizing the cross-entropy loss. In this framework, given sufficient training data, a sufficiently expressive architecture, and effective optimization, the model’s predictions are expected to converge to the true conditional distribution, as follows:
$p(y|\mathbf{x})=\frac{\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y))}{\sum_{y^{\prime}}\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y^{\prime}))}.$ (7)
On the other hand, for the proposed latent variable model described in Eq. 1), the true conditional distribution $p({y}|\mathbf{x})$ can be derived using Bayes’ rule:
$p({y}|\mathbf{x})=\sum\nolimits_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}).$ (8)
@@
$\frac{\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y))}{\sum_{y^{\prime}}\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y^{\prime}))}=\sum\nolimits_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x}).$ (9)
Taking the logarithm on both sides of Eq. (9), we obtain:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log{Z(\mathbf{x})}=\log\big{(}\sum\nolimits_{c}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})\big{)}$ (10)
where ${Z(\mathbf{x})}=\sum_{y^{\prime}}\exp(\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y^{\prime}))$ represents the normalization constant. This transformation provides a direct link between the representations $\mathbf{f}(\mathbf{x})$ learned by next-token prediction (Eq. (7)) and the true posterior distribution $p(\mathbf{c}|\mathbf{x})$.
Now, let us focus on the right of Eq. (10) , we can obtain that:
$\log\big{(}\sum\nolimits_{c}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})\big{)}$ $=\log p(y|\mathbf{x}),$ (11)
$=\mathbb{E}_{p(\mathbf{c}|y)}[\log p(y|\mathbf{x})],$ (12)
$=\mathbb{E}_{p(\mathbf{c}|y)}[\log\big{(}p(\mathbf{c}|y,\mathbf{x})\frac{p(y|\mathbf{x})}{p(\mathbf{c}|y,\mathbf{x})}\big{)}],$ (13)
$=\mathbb{E}_{p(\mathbf{c}|y)}[\log{p(y,\mathbf{c}|\mathbf{x})}]-\mathbb{E}_{p(\mathbf{c}|y)}[\log{p(\mathbf{c}|y,\mathbf{x})}],$ (14)
$=\mathbb{E}_{p(\mathbf{c}|y)}[\log{p(\mathbf{c}|\mathbf{x})}]+\mathbb{E}_{p(\mathbf{c}|y)}[\log{\underbrace{p(y|\mathbf{c},\mathbf{x})}_{p(y|\mathbf{c})}}]-\mathbb{E}_{p(\mathbf{c}|y)}[\log{p(\mathbf{c}|y,\mathbf{x})}],$ (15)
where $p(y|\mathbf{c},\mathbf{x})=p(y|\mathbf{c})$, due to the conditional independence of $y$ and $\mathbf{x}$ given $\mathbf{c}$, as defined in the proposed latent variable model.
@@
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log{Z(\mathbf{x})}=\mathbb{E}_{p(\mathbf{c}|y)}[\log{p(\mathbf{c}|\mathbf{x})}]+\underbrace{\mathbb{E}_{p(\mathbf{c}|y)}[\log{{p(y|\mathbf{c})}}}_{b_{y}}]-\mathbb{E}_{p(\mathbf{c}|y)}[\log{p(\mathbf{c}|y,\mathbf{x})}].$ (16)
Define $\mathbb{E}_{p(\mathbf{c}|y)}[\log{{p(y|\mathbf{c})}}]=b_{y}$, and define a vector $[p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}$ as the vector constructed from the probabilities of all possible values of $\mathbf{c}$, conditional on $\mathbf{x}$, where ${\mathbf{c}_{i}\in\mathcal{C}}$ and $\mathcal{C}$ is the set of all possible values that $\mathbf{c}$ can take. As a result, $\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}=\sum_{\mathbf{c}}{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}=[{p}(\mathbf{c}=\mathbf{c}_{i}|y)]_{\mathbf{c}_{i}}{[\log p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}}$, and similarly $\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|y,\mathbf{x})}=[{p(\mathbf{c}=\mathbf{c}_{i}|y)}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|y,\mathbf{x})}]_{\mathbf{c}_{i}}$. Then we can re-write Eq. (16) as follows:
@@
Now, we apply the diversity condition outlined in Section 3, which asserts that there exist $\ell+1$ values of $y$, i.e., $y_{0}$, $y_{1}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})]^{T}_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]^{T}_{\mathbf{c}_{i}},...,[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})]^{T}_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]^{T}_{\mathbf{c}_{i}}\big{)}$ of size $\ell\times\ell$ is invertible, where $\ell$ is the number of all possible values of $\mathbf{c}$. In this context, for $y=0$, we have:
@@
$\big{(}\mathbf{f}_{y}(y=y_{1})^{T}-\mathbf{f}_{y}(y=y_{0})^{T}\big{)}\mathbf{f}_{x}(\mathbf{x})=$ $\big{(}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})}-{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]^{T}_{\mathbf{c}_{i}}\big{)}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$
@@
According to the diversity condition, where $y$ can take $\ell+1$ values, we can obtain a total of $\ell$ equations similar to Eq. (21). Collecting all of these equations, we have:
@@
$=\underbrace{\big{(}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})}-{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]_{\mathbf{c}_{i}},...,[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})}-{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]_{\mathbf{c}_{i}}\big{)}^{T}}_{\mathbf{L}}\times$
@@
We first focus on the term $\mathbf{b}_{y}$ on the right-hand side of Eq. (24). Note that $b_{y_{i}}=\mathbb{E}_{p(\mathbf{c}\mid y_{i})}[\cdot]$ is a constant with respect to $\mathbf{c}$, as the expectation integrates over all possible values of $\mathbf{c}$. As a result, the entire term $(\mathbf{\hat{L}}^{T})^{-1}\mathbf{b}_{y}$, denoted as $\mathbf{b}$, on the right-hand side of Eq. (24) is also a constant.
We then focus on the term $\mathbf{{h}}_{y}$ in the right of Eq. (24), i.e., $\mathbf{{h}}_{y}$, from the viewpoints of when the mapping $\mathbf{g}$ from latent space, i.e., $\mathbf{c}$, to observed space, i.e., $\mathbf{x}$ and $y$, is invertible and approximately invertible, respectively.
Invertible
When the mapping $\mathbf{g}$ from latent space to observed space is invertible, meaning that for
@@
we have $\epsilon=0$. Then, for $\mathbf{h}_{y}$, we analyze each component, i.e., $h_{y_{i}}-h_{y_{0}}$, where
@@
which implies that the posterior is concentrated at a single point $\mathbf{c}^{*}$, satisfying $\log p(\mathbf{c}^{*}\mid\mathbf{x},y)=0$. In this case, we have
@@
Approximately Invertible
When the mapping $\mathbf{g}$ is approximately invertible, i.e., $\epsilon\rightarrow 0$, for $\mathbf{h}_{y}$, we analyze the difference:
$h_{y_{i}}-h_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(\mathbf{c}\mid\mathbf{x},y_{i})\right]-\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(\mathbf{c}\mid\mathbf{x},y_{0})\right].$ (29)
@@
$p(\mathbf{c}\mid\mathbf{x},y)=(1-\epsilon)\cdot\delta_{\mathbf{c}^{*}}+\epsilon\cdot q(\mathbf{c}),$ (30)
@@
$\mathbb{E}_{p(\mathbf{c}\mid y)}\left[\log p(\mathbf{c}\mid\mathbf{x},y)\right]=p(\mathbf{c}^{*}\mid y)\log(1-\epsilon)+\sum_{\mathbf{c}\neq\mathbf{c}^{*}}p(\mathbf{c}\mid y)\log p(\mathbf{c}\mid\mathbf{x},y).$ (31)
@@
$\mathbb{E}_{p(\mathbf{c}\mid y)}\left[\log p(\mathbf{c}\mid\mathbf{x},y)\right]\approx\log\epsilon,$ (32)